One number is 3 times another. If the sum of their reciprocals is 1/5 , find the two numbers.
Answer by stanbon(48558)

One number is 3 times another. If the sum of their reciprocals is 1/5 , find the two numbers.
-------
One number is "x" ; the other is "3x"
EQUATION:
1/x + 1/3x = 1/5
Multiply thru by 15x to get:
15 +5 = 3x
20 = 3x
x=(20/3) ----this is one of the numbers
3x= 20 ----this is the other number
==============
Cheers,
Stan H.